# Bayesian Model Parameter Learning in Linear Inverse Problems: Application in EEG Focal Source Imaging

### v1

@@ 3.4 Inference of the Model Parameter
3.4.1 Conditional Gaussian Approximation
In Bayesian framework, we expect to obtain model parameter estimates by maximizing the conditional probability distribution
$\hat{\sigma}:=\max_{\sigma}\pi(\sigma|\alpha).$ (18)
In particular, the conditional probability distribution $\pi(\sigma|{\alpha})$ can be expressed as
$\pi(\sigma|\alpha)=\int\pi(\sigma,x|\alpha)\;\mathrm{d}x,$ (19)
where $\sigma$ and $x$ are mutually independent, and $\pi(\sigma)$ can be modeled as a Gaussian distribution with mean $\sigma_{*}$ and variance $\gamma_{\sigma}$. Even when both $\pi(\sigma)$ and $\pi_{x}(.)$ are modeled as Gaussian, the relationship between $x$, $\sigma$ and $\alpha$ can be complicated and cause difficulties for Bayesian inference.
A simple approach is to approximate $\pi(\sigma|\alpha)\propto\pi(\sigma,\alpha)$ and rely on Monte Carlo simulations [37]. In particular, a joint Gaussian distribution for the pair ($\alpha,\sigma$) can be a good approximation for small perturbations $\sigma-\sigma_{*}$. In other words, a first order Taylor approximation with only the linear term in (11) can be sufficient to describe the modeling error. Therefore, given an estimate for the error coefficients $\hat{\alpha}$, the mean of $\pi(\sigma|\hat{\alpha})$ is given by
@@ 3.4.1 Conditional Gaussian Approximation
where $\sigma_{*}$ is the mean value of the postulated model parameter distribution, $\Gamma_{\sigma\alpha}$ is the cross-covariance between $\sigma$ and $\alpha$ (estimated using samples $\alpha^{(s)}$ and $\sigma^{(s)}$) and $\Gamma_{\alpha}=\mathrm{diag}\{\lambda_{1},\ldots,\lambda_{p}\}$.
As we shall see later (in Section 4.2.2), in some cases the CG-based parameter estimation can be accompanied with an iterative algorithm in order to improve the estimates.
3.4.2 Gaussian Process
@@
In more general cases, variables $(x,\alpha)$ and $\sigma$ can be non-linearly related, and therefore we should consider $\sigma=f(\alpha,x)$, where $f$ is an unknown function. To infer $\sigma$, we can model $f$ to be distributed as a Gaussian process, $f\sim\mathcal{GP}(m,K)$, with mean function $m(\alpha,x)=\mathbb{E}[f(\alpha,x)]$ and covariance (or kernel) function $K_{(\alpha,x)}=\mathbb{E}[(f(\alpha,x))-m(\alpha,x))((f(\alpha^{\prime},x^{\prime}))-m(\alpha^{\prime},x^{\prime}))^{\mathrm{T}}]$ [23]. In practice, based on training input-output data, $\mathcal{D}=\{x^{(s)},\alpha^{(s)}\}_{s=1}^{N_{s}}$ and $f_{\mathcal{D}}=\{\sigma^{(s)}\}_{s=1}^{N_{S}}$, respectively, we can infer a new model parameter value $\sigma$ given a new input estimated from (16). Thus, we can predict $\hat{\sigma}$ given $\mathcal{\hat{D}}=(\hat{\alpha},\hat{x})$ from (17) and the training data $\mathcal{D}$. The predictive distribution is $\pi(f_{\mathcal{\hat{D}}}|\mathcal{\hat{D}},\mathcal{D},f_{\mathcal{D}})\sim\mathcal{N}(\hat{f}_{\mathcal{\hat{D}|\mathcal{D}}},K_{\mathcal{\hat{D}}|\mathcal{D}})$ where
$\hat{f}_{\mathcal{\hat{D}|\mathcal{D}}}=m(\hat{D})+K_{\mathcal{\hat{D}},\mathcal{{D}}}K_{\mathcal{D}}^{-1}(f_{\mathcal{D}}-m(\mathcal{D}))$ (21)
@@
is the conditional mean based on the training set, $K_{\mathcal{\hat{D}}|\mathcal{D}}=K_{\mathcal{\hat{D}}}-K_{\mathcal{\hat{D}},\mathcal{{D}}}K_{\mathcal{D}}^{-1}K_{\mathcal{\hat{D}},\mathcal{{D}}}^{\mathrm{T}}$ is the conditional variance, $K_{\mathcal{\hat{D}},\mathcal{{D}}}$ is a vector of covariances between every training case and $(\hat{\alpha},\hat{x})$, $K_{\mathcal{D}}$ is a matrix with the training data set covariance, and $K_{\mathcal{\hat{D}}}$ is the variance of $(\hat{\alpha},\hat{x})$. The choice of the mean function and kernel can be determined based on available (sample) data and properties (e.g. physics) of the inverse problem in question.
4 Application in EEG Source Imaging
@@
4.1 Simultaneous Approximation Error and Source Estimation using Model-based Learning and Dipole Scan
@@
In this work, we apply the described Bayesian framework for the reconstruction of the unknown skull conductivity (model parameter) in the EEG source imaging problem alongside the source activity. Here, we consider the distributed source modeling [50] where the primary unknown $x\in\mathbb{R}^{3n}$ is the distributed (electrical) current dipole source configuration (or field) in a source space that consists of $n$ discrete locations, $\left|x_{i}\right|=\sqrt{\mathrm{x}_{i1}^{2}+\mathrm{x}_{i2}^{2}+\mathrm{x}_{i3}^{2}}$ is the amplitude of the source at location $i$, and $(\mathrm{x}_{i1},\mathrm{x}_{i2},\mathrm{x}_{i3})$ are the components of $x_{i}$ along the coordinate axes. The observation model is linear, as Equation (1), where $v\in\mathbb{R}^{m}$ are the EEG recordings, and the system matrix $A(\sigma)\in\mathbb{R}^{m\times 3n}$ is called the leadfield matrix, which depends on the model parameter, electrical conductivity of the skull $\sigma\in\mathbb{R}^{+}$.
@@
To obtain estimates for the source and approximation error, we first compute the maximum a posteriori (MAP) estimates of the posterior (16). Here, we utilize the dipole scan solver which is a very commonly used approach in EEG studies [51]. In the dipole scan algorithm, the main assumption is that only a single dipole source is active at a time.
Hence, the EEG observation model can be written as
$v=A^{i}_{0}x_{i}+W_{i}\alpha_{i}+\varepsilon_{i*}+\varepsilon_{i}^{\prime\prime}+e.$ (22)
@@
The subscript $i$ denotes a specific dipole location and the superscript in $A^{i}(\sigma)$ denotes the $i1$, $i2$, and $i3$ columns of the leadfield matrix. The approximation error covariance matrix related to the dipole at location $i$ is denoted by $\Gamma_{\varepsilon_{i}}=\mathbb{E}[(\varepsilon_{i}-\varepsilon_{i*})(\varepsilon_{i}-\varepsilon_{i*})^{\mathrm{T}}]$ where $\varepsilon_{i*}$ is the mean approximation error at location $i$. Then, the primary error is $\varepsilon^{\prime}_{i}=W_{i}\alpha_{i}$, where $W_{i}$ are the first $p$ eigenvectors corresponding to the highest eigenvalues obtained from the eigenvalue decomposition of $\Gamma_{\varepsilon_{i}}$.
Based on the model (22), we can compute a MAP estimate of the pair $(x_{i},\alpha_{i})$ at each location $i$, and then select as a solution the pair that maximizes the posterior (16). Therefore, given location $i$, we solve the linear system
@@
$\begin{bmatrix}\hat{x}_{i}\\ \hat{\alpha}_{i}\end{bmatrix}:=\begin{bmatrix}L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}A_{0}^{i}&L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}W_{i}\\ \mathbf{0}&L_{\alpha_{i}}\end{bmatrix}^{-1}\begin{bmatrix}L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}(v-\varepsilon_{i*}-e_{*})\\ 0\end{bmatrix},\;\;\;\mbox{for }i=1,\ldots,n,$ (23)
where $n$ is the total number of source locations, $\hat{x}_{i}$ is the dipole source reconstructed at location $i$, $L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}$ comes the Cholesky factorization of $(\Gamma_{\varepsilon_{i}^{{}^{\prime\prime}}}+\Gamma_{e})^{-1}$, $W_{i}\in\mathbb{R}^{m\times p}$ are the eigenvectors corresponding to the $p$ largest eigenvalues obtained from the eigenvalue decomposition of the error covariance $\Gamma_{\varepsilon_{i}}\in\mathbb{R}^{m\times m}$, $\alpha_{i}\in\mathbb{R}^{p}$ are the error coefficients at this location, and $L_{\alpha_{i}}=\frac{1}{2}\mathrm{diag}\{\lambda_{i1}^{1/2},...,\lambda_{ip}^{1/2}\},$ where $\lambda_{ip}$ are the $p$ largest eigenvalues of the covariance $\Gamma_{\varepsilon_{i}}$. The solution is the pair
$(\hat{x}_{l},\hat{\alpha}_{l})\leftarrow\min_{i=1:n}\{\|L_{\varepsilon^{{}^{\prime\prime}}_{i}+e+\delta_{i}}(v-A_{0}^{i}\hat{x}_{i}-\varepsilon_{i*}-W_{i}\hat{\alpha}_{i}-e_{*})\|_{2}^{2}+\|L_{\alpha_{i}}\hat{\alpha}_{i}\|_{2}^{2}\},$ (24)
where index $l$ refers to the location that gives the lowest value for the functional (24), and the corresponding approximation error is given by $\hat{\varepsilon}^{\prime}_{l}=W_{l}\hat{\alpha}_{l}$. In the following subsections, we analyze the strategies for estimating $\sigma$ given the pair $(\hat{x}_{l},\hat{\alpha}_{l})$.

### v2

@@ 3.3 Inference of the Model Parameter
3.3.1 Variable Interactions
The inference of the model parameter $\sigma$ from the estimated error (as in [43])
$\hat{\varepsilon}\approx W^{p}\hat{\alpha}+\varepsilon_{*}$ (12)
is feasible if the basis functions $W^{p}$ primarily capture the error induced by the unknown model parameter $\sigma$. However, this is not always the case, and in many situations, the estimated signal $\hat{x}$ must also be considered when estimating $\sigma$. Without loss of generality in the following analysis we consider that $\sigma>0$ is a scalar.
In particular, even though $x$ and $\sigma$ are independent variables, $x$ can still influence $\sigma$ through its interaction with $\alpha$. This implies that we are dealing with an interaction effect rather than a direct causal relationship. To better understand how the linear and nonlinear effects of $\sigma$ and $x$ are incorporated into the basis functions $W^{p}$ and, consequently, affect the estimation of $\alpha$ (which we use to infer $\sigma$), we express the error covariance semi-analytically using a Taylor expansion of the error (6).
Expanding around the mean $\sigma_{*}$, we obtain:
$\varepsilon(x,\sigma)=A(\sigma_{*})x-A_{0}x+\sum_{k}J_{k}(\sigma_{*})z_{k}(\sigma)x,$ (13)
where $J_{k}=\frac{1}{k!}\frac{\partial^{k}A(\sigma)}{\partial\sigma^{k}}\Big{|}_{\sigma=\sigma_{*}}\in\mathbb{R}^{m\times n}$ and $z_{k}(\sigma)=(\sigma-\sigma_{*})^{k}$.
For simplicity, we assume $\sigma_{*}=\sigma_{0}$, and the error covariance can be expressed as:
$\Gamma_{\varepsilon}=\sum_{k}J_{k}\,\operatorname{cov}[xz_{k}]\,J_{k}^{\mathrm{T}}+\sum_{k\neq j}J_{k}\,\operatorname{cov}[x{z_{k}},xz_{j}^{\mathrm{T}}]\,J_{j}^{\mathrm{T}}.$ (14)
Since $z_{k}$ and $x$ are mutually independent, it follows from [55] that the covariance matrix of their product is given by:
$\operatorname{cov}[xz_{k}]=\operatorname{cov}[x]\operatorname{var}[z_{k}]+\operatorname{cov}[x]\mathbb{E}[z_{k}]^{2}+\operatorname{var}[z_{k}]\mathbb{E}[x]\mathbb{E}[x]^{\mathrm{T}}.$ (15)
Furthermore, when the mean $\mathbb{E}[x]=0$ and $x$ has isotropic covariance, i.e. $\mathrm{cov}[x]=\gamma I_{n}$, we get
$\Gamma_{\varepsilon}=\gamma\sum_{k}J_{k}\left(\operatorname{var}[z_{k}]+\mathbb{E}[z_{k}]^{2}\right)J_{k}^{\mathrm{T}}+\gamma\sum_{k\neq j}J_{k}\mathbb{E}[z_{k}z_{j}^{\mathrm{T}}]J_{j}^{\mathrm{T}}.$ (16)
This expression for the error covariance suggests that $x$ has primarily a scaling effect on the error covariance. Furthermore, we need to determine whether matrix $J_{k}$ implicitly encodes information about $x$, potentially arising from the modeling process (e.g., linearization of a nonlinear problem). Overall, we expect the basis functions $W^{p}$ (obtained via the eigenvalue decomposition of $\Gamma_{\varepsilon}$) to capture error contributions depending both from the unknown model parameter $\sigma$ and the primary signal $x$ [^2] If $\sigma\,\in\mathbb{R}^{N_{\sigma}}$ where $N_{\sigma}>1$, we can obtain a similar expression, i.e. $\Gamma_{\varepsilon}=\gamma\sum_{k,j}J_{k}\left(\operatorname{cov}[z_{k},z_{j}]\otimes I_{n}\right)J_{j}^{\top}.$, where $J_{k}\,\in\,\mathbb{R}^{m\times(n\cdot N_{\sigma})}$ and $\otimes$ is the Kronecker product..
3.3.2 Conditional Gaussian (CG) Approximation
In the Bayesian framework, we expect to obtain model parameter estimates by maximizing the conditional probability distribution $\hat{\sigma}:=\max_{\sigma}\pi(\sigma|\alpha)$, however this is not always tractable. Instead, a simple approach is to approximate $\pi(\sigma|\alpha)\propto\pi(\sigma,\alpha)$ and rely on Monte Carlo simulations [43]. In particular, a joint Gaussian distribution for the pair ($\alpha,\sigma$) can be a good approximation for small differences $\sigma-\sigma_{*}$. In other words, first order Taylor approximation with only the linear term in (13) is sufficient to describe the modeling error. Therefore, given an estimate for the error coefficients $\hat{\alpha}$, the mean of $\pi(\sigma|\hat{\alpha})$ is given by
@@ 3.3.2 Conditional Gaussian (CG) Approximation
where $\sigma_{*}$ denotes the mean value of the postulated model parameter distribution, the cross-covariance $\Gamma_{\sigma\alpha}$ is estimated using samples of $\alpha^{(s)}$ and $\sigma^{(s)}$, and $\Gamma_{\alpha}=\mathrm{diag}\{\lambda_{1},\ldots,\lambda_{p}\}$. As we shall see later (in Section 4.2.2), in some cases the CG-based parameter estimation can be used to initialize a nonlinear iterative algorithm to improve the model parameter estimates.
3.3.3 Gaussian Process (GP)
@@
In more general cases, where $\alpha=g(x,\sigma)$, we can employ statistical techniques to learn the inverse function, $\sigma=f(\alpha,x)$, in order to predict $\sigma$. As explained earlier (in Sections 3.1 and 3.3.1), this expression does not mean that $x$ and $\sigma$ are dependent; it simply means that given $\alpha$, we may be able to refine our estimate of $\sigma$ by knowing $x$, as the value of $\alpha$ depends on both $\sigma$ and $x$. To infer $\sigma$, we can model $f$ to be distributed as a Gaussian process, $f\sim\mathcal{GP}(m,K)$, with mean function $m(\alpha,x)=\mathbb{E}[f(\alpha,x)]$ and covariance (or kernel) function $K_{(\alpha,x)}=\mathbb{E}[(f(\alpha,x))-m(\alpha,x))((f(\alpha^{\prime},x^{\prime}))-m(\alpha^{\prime},x^{\prime}))^{\mathrm{T}}]$ [23]. In practice, based on training input-output data, $\mathcal{D}=\{x^{(s)},\alpha^{(s)}\}_{s=1}^{N_{s}}$ and $f_{\mathcal{D}}=\{\sigma^{(s)}\}_{s=1}^{N_{S}}$ respectively, we can infer a new model parameter value $\hat{\sigma}$ given a new input estimated from (11). Thus, we can predict $\hat{\sigma}$ given $\mathcal{\hat{D}}=(\hat{x},\hat{\alpha})$ and the training data $\mathcal{D}$.
The predictive distribution is $\pi(f_{\mathcal{\hat{D}}}|\mathcal{\hat{D}},\mathcal{D},f_{\mathcal{D}})\sim\mathcal{N}(\hat{f}_{\mathcal{\hat{D}|\mathcal{D}}},K_{\mathcal{\hat{D}}|\mathcal{D}})$. The conditional mean based on the training set is
$\hat{f}_{\mathcal{\hat{D}|\mathcal{D}}}=m(\hat{D})+K_{\mathcal{\hat{D}},\mathcal{{D}}}K_{\mathcal{D}}^{-1}(f_{\mathcal{D}}-m(\mathcal{D})),$ (18)
@@
where $K_{\mathcal{\hat{D}}|\mathcal{D}}=K_{\mathcal{\hat{D}}}-K_{\mathcal{\hat{D}},\mathcal{{D}}}K_{\mathcal{D}}^{-1}K_{\mathcal{\hat{D}},\mathcal{{D}}}^{\mathrm{T}}$ is the conditional variance, $K_{\mathcal{\hat{D}},\mathcal{{D}}}$ is a vector of covariances between every training case and $(\hat{\alpha},\hat{x})$, $K_{\mathcal{D}}$ is a matrix with the training data set covariance, and $K_{\mathcal{\hat{D}}}$ is the variance of $(\hat{\alpha},\hat{x})$. The choice of the mean function and kernel can be determined based on available (sample) data and properties (e.g. physics) of the inverse problem in question.
4 Dipole Source Imaging and Skull Conductivity Estimation using Somatosensory EEG Data
In this section, we present a framework for joint estimation of brain source activity and skull conductivity using EEG responses. Building on the previously presented Bayesian approach, we present a modified single-dipole scanning technique to estimate single sources and approximate modeling errors. Then, we describe how the modeling error, which is dependent on the skull conductivity, can be exploited to calibrate the skull conductivity value. To achieve this, we use three strategies, a conditional Gaussian (CG) approximation, an iterative algorithm that can refine the CG-result (CG+Iter.), and finally a Gaussian process (GP) -based regression. In Section 5, we demonstrate the proposed approaches with simulated EEG recordings generated by single-dipole sources in the somatosensory cortex of the brain, as this kind of EEG (and MEG) datasets have been shown to be suitable for skull conductivity calibration [33, 34, 35, 37, 38].
@@ 4 Dipole Source Imaging and Skull Conductivity Estimation using Somatosensory EEG Data
4.1 Simultaneous Approximation Error and Source Estimation using Model-based Learning and Single Dipole Scanning
@@
In this work, we apply the described Bayesian framework for the reconstruction of the unknown skull conductivity (model parameter) in the EEG source imaging problem alongside the source activity. Here, we consider the distributed source modeling [56, 57] where the primary unknown $x\in\mathbb{R}^{3n}$ is the distributed (electrical) current dipole source configuration (or field) in a source space that consists of $n$ discrete locations, $\left|x_{i}\right|=\sqrt{\mathrm{x}_{i1}^{2}+\mathrm{x}_{i2}^{2}+\mathrm{x}_{i3}^{2}}$ is the amplitude of the source at location $i$, and $(\mathrm{x}_{i1},\mathrm{x}_{i2},\mathrm{x}_{i3})$ are the components of $x_{i}$ along the coordinate axes. The observation model is linear, as Equation (1), where $v\in\mathbb{R}^{m}$ are the EEG recordings, and the system matrix $A(\sigma)\in\mathbb{R}^{m\times 3n}$ is called the leadfield matrix, which depends on the model parameter, electrical conductivity of the skull $\sigma\in\mathbb{R}^{+}$.
@@
To obtain estimates for the source and the approximation error, we first compute the maximum a posteriori (MAP) estimates of the posterior (11). Since, in the numerical experiments (Section 5), we use simulated EEG recordings generated by single dipole sources within a restricted region (namely, the somatosensory cortex of the brain), we employ the dipole scanning solver that is widely used in EEG studies [58, 59, 60, 61]. In the dipole scanning algorithm, the main assumption is that only a single dipole source $x_{i}$ is active at a time.
In Bayesian perspective, the standard dipole scanning algorithm solves $\hat{x}_{i}=\arg\max_{x_{i}}\pi(v|x_{i})$, where $\pi(v|x_{i})$ is the likelihood considering a single source at location $i$, and the solution is the estimate $\hat{x}_{i}$ for which the likelihood $\pi(v|\hat{x}_{i})$ is maximized.
For the proposed framework, the approximate likelihood of the dipole scanning algorithm is
$\tilde{\pi}(v|x_{i},\alpha_{i})\propto\exp{\frac{1}{2}\left((v-A^{i}_{0}x_{i}+W^{p}_{i}\alpha_{i}+\varepsilon_{i*})^{\mathrm{T}}\Gamma_{e+\varepsilon_{i}^{\prime\prime}}^{-1}(v-A^{i}_{0}x_{i}+W^{p}_{i}\alpha_{i}+\varepsilon_{i*})\right)}.$ (19)
@@
Here the subscript $i$ denotes a specific dipole location, the superscript in $A^{i}(\sigma)$ denotes the $i1$, $i2$, and $i3$ columns of the leadfield matrix and $\alpha_{i}\,\in\mathbb{R}^{p}$ are the error coefficients encoding the modelling error related to location $i$. In particular, the error related to location $i$ is $\varepsilon_{i}(\sigma)=A^{i}(\sigma)x_{i}-A_{0}^{i}x_{i}\,\in\mathbb{R}^{m}$ and its representation using basis functions is $\varepsilon_{i}(\sigma)=W^{p}_{i}\alpha_{i}+\varepsilon_{i*}+\varepsilon_{i}^{\prime\prime}$. The basis functions $W^{p}_{i}\,\in\mathbb{R}^{m\times p}$ are the $p$ top eigenvectors from the eigenvalue decomposition of the approximation error covariance matrix related to a source at location $i$ which is denoted by $\Gamma_{\varepsilon_{i}}=\mathbb{E}[(\varepsilon_{i}-\varepsilon_{i*})(\varepsilon_{i}-\varepsilon_{i*})^{\mathrm{T}}]\,\in\mathbb{R}^{m\times m}$ where $\varepsilon_{i*}\,\in\mathbb{R}^{m}$ is the mean approximation error at location $i$.
Therefore, we can compute a MAP estimate of the pair $(\hat{x}_{i},\hat{\alpha}_{i})$ at each location $i$, and then select as a solution the pair that results in the maximum posterior $\pi({x}_{i},{\alpha}_{i}|v)\propto\pi(v|{x}_{i},{\alpha}_{i})\pi(x_{i},{\alpha}_{i})$. In particular, considering uniform prior for $\pi(x_{i})$ and $\pi(\alpha_{i})\sim\mathcal{N}(0,\Gamma_{\alpha_{i}})$ where $\Gamma_{\alpha_{i}}=\mathrm{diag}\{\lambda_{i1},\ldots,\lambda_{ip}\}$, we have the prior $\pi(x_{i},\alpha_{i})\propto\pi(\alpha_{i})$. Then, the dipole scanning algorithm solves for $i=1,\ldots,n$
$(\hat{x}_{i},\hat{\alpha}_{i})=\arg\max_{x_{i},\alpha_{i}}\pi(v|x_{i},\alpha_{i})\,\pi(\alpha_{i}).$ (20)
The final solution is
$(\hat{x}_{l},\hat{\alpha}_{l})\leftarrow\max_{i=1,\dots,n}\pi(v|\hat{x}_{i},\hat{\alpha}_{i})\,\pi(\hat{\alpha}_{i}),$ (21)
where index $l$ refers to the location that gives the maximum posterior. From (20), (19) and for the given prior, we have that
@@
$(\hat{x}_{i},\hat{\alpha}_{i})=\arg\max_{x_{i},\alpha_{i}}\exp{\left(-\frac{1}{2}\left\|\begin{bmatrix}L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}A_{0}^{i}&L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}W^{p}_{i}\\ \mathbf{0}&L_{\alpha_{i}}\end{bmatrix}\begin{bmatrix}x_{i}\\ \alpha_{i}\end{bmatrix}-\begin{bmatrix}L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}(v-\varepsilon_{i*}-e_{*})\\ 0\end{bmatrix}\right\|_{2}^{2}\right)},$ (22)
where $\|\cdot\|_{2}^{2}$ is the $\ell_{2}$-norm, $L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}$ comes from the Cholesky factorization of $\Gamma_{\varepsilon_{i}^{\prime\prime}+e}^{-1}=(\Gamma_{\varepsilon_{i}^{{}^{\prime\prime}}}+\Gamma_{e})^{-1}$ and $L_{\alpha_{i}}=\frac{1}{2}\mathrm{diag}\{\lambda_{i1}^{-1/2},...,\lambda_{ip}^{-1/2}\}$ where $\Gamma_{\alpha_{i}}^{-1}=L_{\alpha_{i}}L_{\alpha_{i}}^{\mathrm{T}}$. Based on the previous, the MAP estimate for location $i$ is
$\begin{bmatrix}\hat{x}_{i}\\ \hat{\alpha}_{i}\end{bmatrix}:=\begin{bmatrix}L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}A_{0}^{i}&L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}W^{p}_{i}\\ \mathbf{0}&L_{\alpha_{i}}\end{bmatrix}^{-1}\begin{bmatrix}L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}(v-\varepsilon_{i*}-e_{*})\\ 0\end{bmatrix}.$ (23)
Finally, the solution is the pair denoted with subscript $l$ that minimizes the functional
$(\hat{x}_{l},\hat{\alpha}_{l})\leftarrow\min_{i=1:n}\{\|L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}(v-A_{0}^{i}\hat{x}_{i}-\varepsilon_{i*}-W^{p}_{i}\hat{\alpha}_{i}-e_{*})\|_{2}^{2}+\|L_{\alpha_{i}}\hat{\alpha}_{i}\|_{2}^{2}\}.$ (24)
Here, the corresponding primary error is given by $\hat{\varepsilon}^{\prime}_{l}=W^{p}_{l}\hat{\alpha}_{l}$.
In the following subsections, we analyze the strategies for estimating $\sigma$ given the pair $(\hat{x}_{l},\hat{\alpha}_{l})$.